OpenStudy (anonymous):

Giving Medal to each Explained question Find the solution to the system of equations given below using elimination by addition. 8x + 8y = 48 4x - 4y = 40 A. x = 8, y = 2 B. x = -8, y = 12 C. x = 11, y = 1 D. x = 8, y = -2

OpenStudy (anonymous):

8x + 8y -2(4x - 4y) = 48-40*2 16y=-32, thus y=-2, sub in -2 to first eqn, 64=8x, x=8

OpenStudy (anonymous):

trolled

OpenStudy (anonymous):

@jojo4eva do u think you could explain this one to me i am so confused

OpenStudy (jojo4eva):

ok basically you want to cancel out one of em so lets see multiply the bottom equation by -2 so it cancels out the ys

OpenStudy (anonymous):

its seems like the answer choice would be D

OpenStudy (jojo4eva):

ya @trollswilerule is correct starr

OpenStudy (anonymous):

ok so it was D it was just confusing m the way @trollswillrule